# Advancing Human Action Recognition with Foundation Models trained on Unlabeled Public Videos

### v3

@@ 7 Ethics
This research adheres to strict ethical guidelines regarding the use of online video data, specifically videos collected from TikTok. All videos were legally accessed via the TikTok Research API, and only those videos where the user granted download permission were included. Our data collection process is in full compliance with the TikTok Research API Terms of Service. As an additional layer of security, we delete the data after training.
We place a high priority on privacy and copyright issues. To protect individual privacy, we have implemented face obfuscation techniques in the video clips included in our paper. Moreover, in line with the terms of service, we do not distribute any TikTok user content directly. Instead, we focus on training a foundation model, which we only use to identify action patterns, not personal biometric features.
Ethical Considerations in Online Video Research: The use of online video data poses unique challenges and ethical considerations, particularly regarding privacy, informed consent, and the potential risks associated with the use of such data. Our study emphasizes the importance of adhering to ethical principles that safeguard participant interests. We have made significant efforts to minimize potential harm by anonymizing the data and not sharing the scraped dataset. Although techniques like pixelating faces can be circumvented and are not foolproof, they represent our commitment to reducing identifiable information.
Furthermore, we advocate for the inclusion of clearer guidelines in the terms of service on video platforms, which would allow users to explicitly exclude their videos from research use. This approach does not replace the need for informed consent but is a step towards enhancing ethical practices in online video research. We aim to foster a transparent and informed discussion about the use of online video data in computational research.
@@ Acknowledgements
This work was made possible by a multitude of supporters and funding sources. We thank the National Science Foundation for their generous support through the CyberTraining grant (Award No. 2118222) under the HI-DSI project. The technical assistance and advanced computing resources provided by the University of Hawaii Information Technology Services–Cyberinfrastructure, partially funded by NSF CC* awards #2201428 and #2232862 (Koa), have been invaluable. The development and acceleration of cloud computing through JetStream2 have significantly contributed to our research; we acknowledge David Y. Hancock, Jeremy Fischer, John Michael Lowe, Winona Snapp-Childs, Marlon Pierce, Suresh Marru, J. Eric Coulter, Matthew Vaughn, Brian Beck, Nirav Merchant, Edwin Skidmore, and Gwen Jacobs for this computational infrastructure (DOI: 10.1145/3437359.3465565). The AWS Cloud Credit for Research program has also played a crucial role in providing computational resources for our work. We are grateful for the funding provided by the National Institutes of Health under the NIH Director’s New Innovator Award (award DP2-EB035858). We used ChatGPT to edit the grammar of our manuscript and to re-phrase sentences that were originally worded unclearly. However, all contents in this manuscript are original ideas and analyses conducted by the authors.
Appendix A Appendix
@@ Appendix B Sample Clips
Figure 3: Example classes from the TikTokActions dataset.
Appendix C Dataset Composition
@@
Appendix D Dataset Duration

### v4

@@ 8 Ethics
This study has been approved under the University of Hawaii Institutional Review Board (IRB) under protocol number 2024-00396. Furthermore, this research adheres to strict ethical guidelines regarding the use of online video data, specifically videos collected from TikTok. All videos were legally accessed via the TikTok Research API, and our data collection process is in full compliance with the TikTok Research API Terms of Service. Importantly, only those videos where the user granted explicit download permission were included in our training process. Equally as importantly, we deleted all of the data after model training as an additional layer of security.
We place a high priority on user privacy. To protect individual privacy, we have implemented face obfuscation techniques in the video clips included in our paper. Moreover, in line with the terms of service, we do not distribute any TikTok user content directly. Instead, we focus on training a foundation model, which we only use to identify action patterns, not personal biometric features.
Genderal Ethical Considerations for Online Video Research: The use of online video data poses unique challenges and ethical considerations, particularly regarding privacy, informed consent, and the potential risks associated with the use of such data. Our study emphasizes the importance of adhering to ethical principles that safeguard participant interests. We have made significant efforts to minimize potential harm by anonymizing the data and not sharing the curated dataset. We recommend that researchers adhere to such privacy protections.
Furthermore, we advocate for the inclusion of clearer guidelines in the terms of service on video platforms, which would allow users of the platforms to explicitly exclude their videos from research use. This approach does not replace the need for informed consent but is a step towards enhancing ethical practices in online video research. We aim to foster a transparent and informed discussion about the use of online video data in computational research.
We also argue that the use of the models plays a critical role in deciding the ethical standards that are required. Predicting human actions such as running, swimming, or playing piano from videos is a common task that is unlikely to lead to harm, especially given how common action recognition computer vision research is. By contrast, if the AI were used to predict sensitive properties about the end user such as mental health diagnoses, income, political party, or anything else along those lines, then the standard for explicit consent and ethical practices would rise to a much higher level [39].
[…]
@@ Acknowledgements
This work was made possible by a multitude of supporters and funding sources. We thank the National Science Foundation for their generous support through the CyberTraining grant (Award No. 2118222) under the HI-DSI project. The technical assistance and advanced computing resources provided by the University of Hawaii Information Technology Services–Cyberinfrastructure, partially funded by NSF CC* awards #2201428 and #2232862 (Koa), have been invaluable. The development and acceleration of cloud computing through JetStream2 have significantly contributed to our research; we acknowledge David Y. Hancock, Jeremy Fischer, John Michael Lowe, Winona Snapp-Childs, Marlon Pierce, Suresh Marru, J. Eric Coulter, Matthew Vaughn, Brian Beck, Nirav Merchant, Edwin Skidmore, and Gwen Jacobs for this computational infrastructure (DOI: 10.1145/3437359.3465565). The AWS Cloud Credit for Research program has also played a crucial role in providing computational resources for our work. We used ChatGPT to edit the grammar of our manuscript and to re-phrase sentences that were originally worded unclearly. However, all contents in this manuscript are original ideas and analyses conducted by the authors.
Appendix A Sample Clips
@@
Figure 5: Example clips from the TikTokActions dataset.
Appendix B Dataset Composition
@@
Appendix C Dataset Duration